Find the amount (future value) of the ordinary annuity. (Round your answer to the nearest cent.) 

$1400/semiannual period for 8 yr at 9%/year compounded semiannually 


Find the amount (future value) of the ordinary annuity. (Round your answer to the nearest cent.) 

$100/month for 14 yr at 15%/year compounded monthly
Answer #1

SOLUTION :


First problem :


Period of compounding : semiannually.

Interest rate per period, r = 9/2 = 4.5% = 0.045

=> 1 + r = 1.045

No. of periods, n = 8 * 2 = 16 sem-iannual.

Annuity , A = $1400 per period.


So,


FV 

= A((1+r)^n - 1)/r

= 1400(1.045^16 - 1)/(0.045)

= 31807.07 ($) 


FV of the given annuity = $31,807.07  (ANSWER).


Second problem :


Period of compounding : monthly.

Interest rate per period, r = 15/12 = 1.25% = 0.0125

=> 1 + r = 1.0125

No. of periods, n = 14 * 1 2 = 168 months

Annuity , A = $100 per period.


So,


FV 

= A((1+r)^n - 1)/r

= 100(1.0125^168 - 1)/(0.0125)

= 56484.50 ($) 


FV of the given annuity = $56,484.50 (ANSWER).



answered by: Tulsiram Garg
Answer #2

A = 1400 t = 8 years rate , x = 99% =0-09 - 2 (compounded semi-an Future value is given by FV = 1400 ((1+0.89) 28_17 = 1400 0.I hope you have got your answer , if you need any further clarification in any step, please let me know through comment.